A loading dose of voriconazole consists of 5 mg for every kilogram of body mass. This dose is administered as two IV infusions separated by 12 hours. The concentration of the solution is 10 mg/ml. If a patient weighs 132 pounds, how many milliliters of voriconazole are administered in a day?
Answer #1
  • Loading dose of voriconazole consists of 5 mg for every kilogram of body mass.
  • This dose is administered as two IV infusions separated by 12 hours.
  • The concentration of the solution is 10 mg/ml.
  • Patient weighs 132 pounds = 59 kg
  • Milliliters of voriconazole administered in a day = 59 ml

Calculation:

for 12 hours = 5 mg * 59 kg = 295 mg

for one day (24 hours) = 295 * 2 = 590 mg

Conversion to milliliters for 10 mg/ml solution = 590/10 = 59 ml